# Astrobites Collaboration Improving Undergraduate Astronomy Students’ Skills with Research Literature via Accessible Summaries: A Case Study with Astrobites-based Lesson Plans

### v2

@@ Abstract
Undergraduate physics and astronomy students are expected to engage with scientific literature as they begin their research careers, but reading comprehension skills are rarely explicitly taught in major courses. We seek to determine the efficacy of lesson plans designed to improve undergraduate astronomy (or related) majors’ perceived ability to engage with research literature by using accessible summaries of current research written by experts in the field. During the 2022-2023 academic year, twelve faculty members incorporated lesson plans using accessible summaries from Astrobites into their undergraduate astronomy major courses, surveyed their students before and after the activities, and participated in follow-up interviews with our research team. Quantitative and qualitative survey data clearly show that students’ perceptions of their abilities with jargon, identifying main takeaways of a paper, conceptual understanding of physics and astronomy, and communicating scientific results all improved with use of the tested lesson plans. Additionally, students show evidence of increased confidence of their abilities within astronomy after exposure to these lessons, and instructors valued a ready-to-use resource to incorporate reading comprehension in their pedagogy. This case study with Astrobites-based lesson plans suggests that incorporating current research in the undergraduate classroom through accessible literature summaries may increase students’ confidence and ability to engage with research literature, as well as their preparation for participation in research and applied careers.
@@ I Introduction
Research experience is an integral part of modern STEM undergraduate education, especially for students who plan to continue on to graduate study in the field, and undergraduates are often highly encouraged to get involved in research early in their careers. Exposure to research during undergraduate education also provides a number of benefits, including helping students build research skills, providing a basis for career planning, and encouraging students to develop scientific mindsets [1]. Exposure to current research in the classroom is also known to have a positive impact on undergraduate learning experiences, as current literature directly shows students what questions they can answer in their future studies and provides real-world motivation for course content [2, 3]. Recent work exploring Course Based Undergraduate Research Experiences (CUREs) in physics and astronomy further supports the positive benefits of including research in the classroom [4, 5, 6, 7, 8].
@@
However, scientific research requires the ability to comprehend the literature of a field, and jargon and niche topics used in scientific literature are often a significant barrier for budding scientists. This barrier can be particularly high for underrepresented students who may not have the support or experience needed to parse dense and jargon-heavy papers [9, 10, 11]. Reading comprehension skills are also rarely explicitly taught in the higher education science classroom, despite the fact that research shows that interventions focused on discipline-based reading and writing skills are beneficial to students–for example, these interventions can empower students to better contextualize their work and engage in scientific discourse, making them more effective, culturally-aware, and well-rounded scientists [12, 13, 14, 15, 16]–and correlate with success in STEM courses [17]. The recent Phys21 report from the Joint Task Force on on Undergraduate Physics Programs from the American Physical Society, National Science Foundation, and American Association of Physics Teachers further emphasized the need for literature and communication skills as preparation for 21st century careers, and the lack of existing preparation for said skills; for example, they state, “Unless they write a senior thesis, undergraduates are also not often called upon to search the literature; read, analyze, evaluate, interpret, and cite technical articles; and make specific use of the scientific and engineering information therein, despite the fact that graduates are likely to be called upon to do so whether they pursue graduate study or enter the workforce” [18]. Reading and writing skills may be neglected due to a number of reasons: lack of priority in an already content-packed curriculum [19], lack of awareness of current education research and methods [20, 21, 22], lack of preparation for teaching these skills [23], and a lack of time and resources to implement new lessons [24, 25]. Despite these challenges, there already exist natural opportunities to explicitly introduce reading and writing skills into physics and astronomy curricula, e.g. via lab courses and CUREs.
@@
Some prior work on reading comprehension in the science classroom has been completed in physics, including the use of metacognition and question formulation for improving reading skills [26, 27]; however, this research is generally limited, hyper-focused on a specific reading strategy, and not involving authentic research literature. The available literature on reading comprehension in astronomy-specific courses is even more limited (e.g. one study on reading for introductory non-major courses [28]), and the unique culture and demographic of each STEM sub-field prevents us from blindly extrapolating from the research of other disciplines, such as biology or chemistry where research comprehension has been more commonly studied with findings that support the importance of these skills (e.g. [29].
@@
However, teaching reading comprehension skills for research literature is only one way to tackle this problem; research literature can also be made more accessible via simplified summaries written by experts in the field [30, 22]. Accordingly, we propose that offering easy-to-implement, readily available plans featuring both reading comprehension skills and accessible summaries of recent literature will be beneficial to student skills and confidence with research literature, as well as the uptake of educators actually implementing these lesson plans. To this end, we have developed four astronomy-focused lesson plans to improve students’ skills and confidence with research literature via hands-on experiences with scientific articles, The concept of using accessible summaries (specifically “science bites”) articles for reading comprehension exercises in the classroom was originally presented in Sanders et al. [31], and practical infrastructure, in the form of sample lesson plans, are provided in Sanders et al. [32]. Building from these works, this investigation aims to to gauge the efficacy of interventions focused on reading comprehension and communication skills involving simplified current literature summaries in achieving the following goal: improving student confidence with and comprehension of research literature.
We use lesson plans from the Astrobites website111astrobites.org as an example of such interventions, making this a case study from which further research can build to more generalizable conclusions. These lesson plans harness existing accessible summaries of research literature hosted on the well-known Astrobites platform supported by the American Astronomical Society. This website provides accessible daily summaries of recent research in the field intended for an undergraduate audience, as well as other professional development and education resources for astronomers. In this work, we distributed the aforementioned lesson plans and oversaw their incorporation into undergraduate astronomy courses during the 2022-2023 academic year. Although existing Astrobites resources can be adapted for a variety of educational levels, this study focuses on the applications of these lesson plans to undergraduates who are enrolled in an astronomy or related major 222Related majors include physics, astrophysics, biology, math, geology, engineering, chemistry, computer science, planetary science, and undeclared majors.
In Section II, we describe the study population, further details on the lesson plans used, and the design of the surveys and other assessments used in this case study. In particular, we seek to evaluate the effects of the lessons on their perceived ability to engage with astronomy literature (particularly with respect to parsing jargon, understanding the main takeaways of a paper, their conceptual understanding/intuition, and their ability to communicate about science) and their perceptions of their broader ability and belonging in the field of astronomy. In Section III, we discuss the effects of these lesson plans on students’ perceived abilities, as well as instructor and student feedback on these lesson plans. Finally, in Section IV, we further explore the impacts of these lesson plans as shown in our results, and suggest future uses and investigations inspired by our findings.
@@ II Methods
II.1 Lesson Plans
The Astrobites lesson plans can be summarized as follows:
1.
Periodic Reading Assignments: Students are assigned Astrobites articles and an associated list of guided questions which evaluate their conceptual understanding and reading comprehension.
2.
Student Research Project: Students use Astrobites articles to carry out a research project (delivered as a paper and/or presentation) on an astronomy topic of their choice.
3.
Student Writing Assignment: Students select a research paper and write an Astrobites-like article summarizing the context, content, and key conclusions.
4.
Student Presentation: Students deliver an in-class presentation based on an Astrobites interview post and paper by the interviewee.
Each lesson plan involves reading about current research topics, whether via an Astrobites article and/or original research literature. More detailed descriptions can be found in Appendix C, and further detail about Astrobites more generally is provided in Appendix A. Each participating instructor corresponded with the research team to implement one or more of the lesson plans of the instructor’s choice in their classroom. All instructors were required to participate in a pre-intervention meeting to ensure knowledge of study protocols (including how to distribute student survey assessments) and a post-intervention meeting to gather information on their experience incorporating the lesson plans in their classes. Research team members only advised the instructor, answered questions and did not have any direct contact with the students themselves. Instructors notified the study team of any deviations from the lesson plan as provided. Some instructors were allowed to implement multiple lessons or repeat a lesson multiple times, although the analysis of “dosing effects” from exposure to multiple lessons is not included in this work.
II.2 Assessment Design & Study Population
Student surveys were administered by participating instructors before and after implementing their chosen lesson plan in the form of a Google Forms survey. The survey contained both Likert-scale quantitative questions and open-ended written response questions, probing student perceptions of ability over six categories: parsing jargon, understanding main takeaways, conceptual understanding/intuition, science communication ability, broad ability within astronomy, and feelings of belonging within astronomy. Since the survey did not directly assess student ability, only their perceptions of their ability, we cannot make strong claims about changes in their actual abilities or skills, only their confidence. Post-intervention surveys also included direct questions about lesson plan itself. Student responses were not anonymous, but identifying information was only used to match pre- and post-survey responses and categorize students to their instructor to note the lesson plan they completed. Additionally, instructors participated in post-lesson interviews to gather qualitative information about the efficacy of the lesson plans and avenues for future improvements. All study protocols and assessments described within were approved by both UCLA and University of Virginia’s Institutional Review Boards (UCLA IRB#22-001473 and UVA IRB#5514). Full text of the assessment questions is available in Appendix B.
202 students from 13 courses for astronomy-related majors taught by twelve faculty instructors across nine institutions participated in this study from Fall 2022 to Winter and Spring 2023. Student participants were undergraduates with an intention to major in astronomy or a related subject, and were either taking the course featuring the Astrobites lesson plan for major or minor credit, or intending to pursue an astronomy-related career. Six of the host institutions (hosting eight of the courses) were R1 universities, while three were liberal arts colleges (hosting five of the courses).
Lesson Plan (LP) 1 was taught in eight courses; LP 2 in two courses; a combination of LPs 1 and 2 in one course; a combination of LPs 2 and 3 in one course; a combination of LPs 2, 3, and 4 in one course. Out of these 202 students, 82 responded to both the pre-lesson and post-lesson surveys; all instructors who completed their participation in the study confirmed distributing the surveys as we instructed, so we suspect the lower response rate in the post-lesson surveys, leading to the smaller sample size, is due to an increase in student responsibilities at the end of the school term when these were often administered. Of the 82 students who participated in both surveys, 51 were part of Lesson Plan 1, 7 students were in LP 2, and 16 students were in LP 4. A further 8 students had a combination of Lesson Plan 1 and 2. No students in both surveys were part of LP 3, neither on itself or in combination with other LPs.
The qualitative and quantitative data collection for this study took place in parallel. Students responded to both types of questions in a single survey. The investigation is a convergent mixed-methods study, wherein qualitative and quantitative data are analyzed separately and then synthesized to provide a broader view of the problem; as stated in Ponce and Pagán-Maldonado [35], “The quantitative approach measures the objective aspects of the problem and the qualitative phase enters the subjective aspects of the problem or the experiences of the participants.” A consistency analysis was also completed between the results of the qualitative and quantitative data collected, as described in Appendix F.
@@ II.2 Assessment Design & Study Population
All questions, both qualitative and quantitative, were designed specifically for this study. Assessments were developed in collaboration with UCLA’s Center for the Integration of Research, Teaching, and Learning (CIRTL), and drew on established assessments from education research literature as inspiration to create the surveys [36, 37, 38, 39]. The assessments were pilot tested with astronomy graduate students, who likely have an experience level intermediate to the undergraduates experiencing the lessons and the instructors administering them, to ensure the question text was interpreted reliably and as expected [40, 41].
II.3 Data Analysis
II.3.1 Quantitative Analysis
The data set for quantitative analysis consists of 82 complete pre- and post-survey responses to the 35 Likert-scale questions listed in Appendix B. We summarize the change in the responses to each question with the 95% confidence interval (‘CI95%’) for the mean of the pre-post survey differences in student responses. Using the statistics package Pingouin [42], we perform a paired t-test for each question by pairing pre- and post-survey responses for individual students. For each question, the p-value was determined using the canonical significance threshold of 0.05 [43]. The effect size is measured by Cohen’s [44], with values above classified as ‘moderate,’ values around as ‘small’, and smaller values as ‘very small.’ The Bayes factor in favor of the alternative hypothesis (‘BF10’) (e.g., Bartoš and Wagenmakers [45]) is computed and describes how many times more likely the observed data are under the alternative hypothesis, ‘H1’: that student perceptions changed. The survey questions are designed to fit into six conceptual categories: jargon, main takeaways, conceptual, communication, ability, and belonging (Table 1). Each question in a category is designed to probe a similar underlying concept; Figure 1 shows the degree to which the questions are related to each other based on student responses. We also analyzed changes within these six categories; the same statistical analyses performed with responses to individual questions were applied to each of the conceptual categories. “Reversed” Likert scale questions were adjusted 333For quantitative questions 2 and 5, low Likert scores (disagreement) represent greater confidence. This is the opposite of the majority of the questions, for which higher Likert scores indicate greater confidence. by reversing the Likert scale for these responses prior to performing any group calculations. We also performed an analysis of second-order correlations, described in Appendix G.
Category | Question numbers in category |
Jargon | 1-5 |
Main Takeaways | 6-11 |
Conceptual | 12-18 |
Communication | 19-24 |
Ability | 25-29 |
Belonging | 30-35 |
II.3.2 Qualitative Analysis
We explored student responses to open-ended survey questions regarding perception of ability (Questions 35 through 39, 41, and 44 in Appendix B) through reflexive thematic analysis [47]. Codes and sub-codes were determined inductively, based on the content of student responses. All responses were read before creating codes, and then responses were re-read several times to refine the initial set of themes/codes. After collective creation of a set of thematic categories by the study team (referred to herein as codes, and summarized in Table 2), five raters (authors BLL, GMD, SH, RL, and IM) independently assigned codes to each response. Codes applied by a majority of raters were included as the final set of identified codes for each response, essentially determining the interpretation of the response by consensus; codes applied by raters were not included, which was relatively uncommon ( of codes were rejected).
We do not compute inter-rater reliability scores here, instead following the philosophy that coding is a flexible and organic process that cannot be separated from the perspectives of the researchers themselves, and there is no single “accurate” coding of a qualitative data set [48, 47]. It is worth noting that some students left questions blank, or provided responses that were illegible or unable to be matched to a code. These responses have all been excluded from the analysis, and as a result, the number of respondents per question varies; this is accounted for by reporting percentages instead of raw counts. Responses to student survey feedback (Questions 40, 42, and 43 in Appendix B) and instructor interviews are read holistically instead of coded and their broad themes summarized in Section III.2.
Qualitative response codes were tallied, and the percentage of responses in each main code category (e.g., not including sub-codes) was compared between pre- and post-lesson surveys. We report the percent change between pre- and post-lesson in each relevant category and the tallies of post-lesson responses in Section III, Figures 3 and 4, respectively.
Survey Question |
Codes |
How do you feel about reading scientific research papers? |
Intimidation (Int) (intimidated, daunting, overwhelmed, confused, not comfortable); Require effort (E) (time-consuming, tedious, read multiple times); Worthwhile (W) (rewarding, useful, necessary); Unprepared (U) (lacking knowledge, not taken enough classes, “not ready yet”); Improved (Imp) (e.g., “better now”); Confidence (Co) (comfortable, good, “like I can do it”); Cause of feelings (C) [jargon (j), figures (f), writing style (s), audience (a)]; Enjoyment (En) (interesting, fun, enjoy) |
What comes to mind when you think of astronomy research? |
Topics (T) (cosmology, astrobiology, stars, galaxies, exoplanets, etc.); Perceptions of Research (P) [difficult (d), interesting (i), fun (f), tedious (t), requires preparation (p) (professors, competence, education)], Skills (S) [coding (c) (computers, plotting, data analysis), observation (o) (telescopes, observing), math (m), reading papers (r), writing (w)], Career desires (C) (e.g. ”want to be a part of it”), Scientific process/discovery (D) |
Do you think you are capable of being an astronomer (or related scientist)? Why or why not? |
Yes (Y); No (N); Maybe (M) (depends, yes and no); Confident (C) (good at it, previous or current experience, have what it takes, success in classes); Insecure (I) (no, maybe, unsure, grades, imposter syndrome, doubt); Passion (P) (life long interest, plans for graduate school); Barriers (B) [learning disorder (d), mental health (h), minority (m)] |
Do you feel like you belong in astronomy? Why or why not? |
Yes (Y); No (N); Maybe (M); Motivation (Mo) (motivation, passion, interest); Experience (E) (classes, research, academic success); Community (C) (welcoming, representation, role models, friends + peers); Toxicity (T) (toxic, unwelcoming, lack of representation); Self-doubt (S) (imposter syndrome, negative comparison to peers, don’t feel smart, lack of academic success, feel overwhelmed) |
What is your main takeaway from this activity? |
Astronomy knowledge (K) (topics, interpreting plots, data); Understanding of field (U) (breadth of topics, better, how papers written, programming, statistics, visuals); Confidence (C) (boost, more capable, more confident, less daunting, more daunting, decreased); Accessibility (A) (more accessible, digestible, resources); Motivation (M) |
How have your feelings about astronomy research changed as a result of this activity, if at all? |
Improved (I) (more capable, less daunting, more confident); Enjoyment (E) (appreciation for sci. comm., more inspired to do astro., more excited to read papers); Meta-knowledge of field (M) [Resources (r) (e.g. better knowledge of resources available), Options (o) (better idea of possibilities), Appreciation (a) (appreciation for what goes into research)]; No change (Nc) |
III Results
III.1 Observed Changes in Student Attitudes
Summaries from the quantitative analysis are shown in Table 3 for the conceptual categories and Table 4 for selected individual questions displaying particularly strong evidence for change: p-value 0.05, the largest effect sizes (measured by Cohen’s d), and moderate to extreme evidence for change (measured by BF10). Histograms and statistical summaries for all of the questions are shown in Figure 6 and Table 5 in Appendix E.
All questions but one showed positive mean of differences, indicating an increase in Likert score, corresponding to an average increase in perceived abilities between the pre- and post-surveys. The exception is Question 26, “I am capable of succeeding in my astronomy courses,” which had a mean difference of zero and nonsignificant p-value. The histogram of pre- and post- responses for this question (Figure 7) demonstrates that the majority of students perceived themselves as capable of succeeding in astronomy courses both before and after the intervention.

### v3

# Astrobites Collaboration Improving Undergraduate Astronomy Students’ Skills with Research Literature via Accessible Summaries: An Exploratory Case Study with Astrobites-based Reading Assignments

@@ Abstract
Undergraduate physics and astronomy students are expected to engage with scientific literature as they begin their research careers, but reading comprehension skills are rarely explicitly taught in major courses. We seek to determine the efficacy of a reading assignment designed to improve undergraduate astronomy (or related) majors’ perceived ability to engage with research literature by using accessible summaries of current research written by experts in the field. During the 2022-2023 academic year, faculty members from six institutions incorporated reading assignments using accessible summaries from Astrobites into their undergraduate astronomy major courses, surveyed their students before and after the activities, and participated in follow-up interviews with our research team. Quantitative and qualitative survey data from 52 students show that students’ perceptions of their abilities with jargon and identifying main takeaways of a paper significantly improved with use of the tested assignment template. Additionally, students report increased confidence of their abilities within astronomy after exposure to these assignments, and instructors valued a ready-to-use resource to incorporate reading comprehension in their pedagogy. This exploratory case study with Astrobites-based assignments suggests that incorporating current research in the undergraduate classroom through accessible literature summaries may increase students’ confidence and ability to engage with research literature, assisting in their preparation for participation in research careers.
@@ I Introduction
Research experience is an integral part of modern STEM undergraduate education, especially for students who plan to continue on to graduate study in STEM, and undergraduates are often highly encouraged to get involved in research early in their careers. Exposure to research during undergraduate education also provides a number of benefits. This includes helping students build research skills, providing a basis for career planning, and encouraging students to develop scientific mindsets [1]. Exposure to current research in the classroom is also known to have a positive impact on undergraduate learning experiences: engaging with current scientific literature illustrates the open questions and active areas of research in a field and provides real-world motivation for course content [2, 3]. Recent work exploring Course Based Undergraduate Research Experiences (CUREs) in physics and astronomy further supports the positive benefits of including research in the classroom, including increased persistence in STEM disciplines, development of research skills, stronger identities as scientists, and more [4, 5, 6, 7, 8].
@@
However, reading comprehension skills are also rarely explicitly taught in the higher education science classroom, despite the fact that research shows that interventions focused on discipline-based reading and writing skills are beneficial to students. These interventions can empower students to better contextualize their work and engage in scientific discourse, making them more effective, culturally-aware, and well-rounded scientists [9, 10, 11, 12, 13]. These interventions also correlate with success in STEM courses [14]. The recent Phys21 report from the Joint Task Force on Undergraduate Physics Programs from the American Physical Society, National Science Foundation, and American Association of Physics Teachers further emphasized the need for literature and communication skills as preparation for 21st century careers and the lack of existing preparation for said skills. For example, they state, “Unless they write a senior thesis, undergraduates are also not often called upon to search the literature; read, analyze, evaluate, interpret, and cite technical articles; and make specific use of the scientific and engineering information therein, despite the fact that graduates are likely to be called upon to do so whether they pursue graduate study or enter the workforce” [15]. Reading and writing skills may be neglected due to a number of reasons related to limitations of the instructor: lack of priority in an already content-packed curriculum [16], lack of awareness of current education research and methods [17, 18, 19], lack of preparation for teaching these skills [20], and a lack of time and resources to implement new lessons [21, 22]. Despite these challenges, there already exist natural opportunities to explicitly introduce reading and writing skills into physics and astronomy curricula, e.g. via lab courses and CUREs.
@@
Some prior work on reading comprehension in the science classroom has been completed in physics, including the use of metacognition (i.e. reflection about thought processes) and students’ question formulation for improving reading skills [23, 24]. However, this research is generally limited, hyper-focused on a specific reading strategy, and not grounded in authentic scientific research literature. The available literature on reading comprehension in astronomy-specific courses is even more limited (e.g. one study on reading for introductory non-major courses [25]). In addition, the unique culture and demographics of each STEM sub-field (i.e. physics being substantially less gender balanced, [26]) prevents us from blindly extrapolating from the research of other disciplines (such as biology or chemistry where research comprehension has been more commonly studied with findings that support the importance of these skills, e.g. [27]), due to the substantial differences in the target population.
Research literature can also be made more accessible via simplified summaries written by experts in the field [28, 19]. Accordingly, offering easy-to-implement, readily available assignments featuring both reading comprehension skills and accessible summaries of recent literature may be beneficial to student skills and confidence with research literature, as well as the uptake of educators actually implementing these assignments.
@@
The concept of using accessible summaries (specifically “science bites”) articles for reading comprehension exercises in the classroom was originally presented in Sanders, Kohler, and Newton 2012 [29] while practical infrastructure, in the form of sample assignments, is provided in Sanders et al. 2017 [30]. Building from these works, this investigation aims to gauge the efficacy of interventions focused on reading comprehension and communication skills involving simplified current literature summaries in achieving the following goal: improving student confidence with and comprehension of research literature. We propose that accessible summaries of research literature can be a powerful step towards improving students’ skills and comfort with primary research papers.
We use a reading comprehension assignment from the Astrobites website111astrobites.org as an example of such interventions, making this an exploratory case study from which further research can build to more generalizable conclusions; in particular, we pursue a convergent mixed methods case study, in which different types of data (qualitative, quantitative) are analyzed in conjunction for a specific sample of astronomy undergraduate students. This reading assignment harnesses existing accessible summaries of research literature hosted on the Astrobites website supported by the American Astronomical Society (further information on Astrobites is available in Appendix A). In this work, we distributed the aforementioned assignment and oversaw its incorporation into undergraduate astronomy courses during the 2022-2023 academic year.
Although existing resources from Astrobites can be adapted for a variety of educational levels, this study focuses on the applications of this assignment to undergraduates who are enrolled in an astronomy or related majors11footnotetext: Related majors include physics, astrophysics, biology, math, geology, engineering, chemistry, computer science, planetary science, and undeclared majors.. We seek to evaluate the effects of the assignment on students’ perceived ability to engage with astronomy literature (particularly with respect to parsing jargon, understanding the main takeaways of a paper, their conceptual understanding/intuition, and their ability to communicate about science) and their perceptions of their broader ability and belonging in the field of astronomy.
In Section II, we detail the reading assignment intervention, the design of the assessments used in this study, and the study population. In Section III, we present the analysis and results of the quantitative (Likert scale) survey data, and in Section IV, we present the analysis and results of the qualitative survey responses from students and interviews with educators. The qualitative and quantitative data are brought together in a consistency analysis in Section V. Finally, in Sections VI and VII, we explore the impacts of this assignment as shown in our results and suggest future uses and investigations inspired by our findings.
@@ II Methods
II.1 Intervention
The Astrobites reading assignment template (referred to as “Lesson Plan Type 1: Periodic Astrobites Reading Assignment” in the Astrobites resources) was originally developed and presented in Sanders et al. [29] and Sanders et al. [30]. In this assignment, students are asked to read an assigned Astrobites article and respond to guided questions that test reading comprehension and conceptual understanding. Students are graded on their responses to the questions, which can be gathered electronically through an online form. The questions are discussed in class to promote greater understanding.
The Astrobites assignment template provides instructions and guidelines for adaptations to instructors, but leaves the specific Astrobite and reading questions to the instructor implementing the lesson. This assignment template has already been in use for over five years by educators across the U.S., and educators in this study voluntarily opted in to adopting the study intervention based on the merits of the materials available. Three other assignments that use Astrobites summaries are available (a research project, writing project, and presentation – see Appendix C for details), but this study is limited to the reading assignment.
II.2 Assessment Design
@@
All questions, both qualitative and quantitative, were designed specifically for this study. Assessments were developed in collaboration with UCLA’s Center for the Integration of Research, Teaching, and Learning (CIRTL), and drew on established assessments from education research literature as inspiration to create the surveys [31, 32, 33, 34]. The survey questions were designed to fit into six conceptual categories of interest to the research team, based on the stated learning objectives of the assignment templates: perception of ability to parse jargon (abbreviated as “jargon”), perception of ability to extract main takeaways (“main takeaways”), perception of ability to understand astronomical concepts (“conceptual”), perception of ability to communicate science (“communication”), perception of general ability in the field (“ability”), and sense of belonging (“belonging”) as summarized in Table 1. Each question in a category is designed to probe a similar underlying concept. The assessments were tested with astronomy graduate students, who likely have an experience level intermediate to the undergraduates experiencing the lessons and the instructors administering them, to ensure the question text was interpreted reliably and as expected [35, 36]; only minor wording revisions were needed as a result of this testing, and revisions were decided based on think-aloud interviews with the testing participants where we asked testing participants to verbalize their thoughts in response to question text [37, 38].
Student surveys were administered by participating instructors before and after implementing the assignment in the form of a Google Forms survey. The survey contained both Likert-scale quantitative questions and open-ended written response questions. The qualitative and quantitative data collection for this study took place in parallel. Students responded to both types of questions in a single survey. The investigation is a convergent mixed-methods study, wherein qualitative and quantitative data are analyzed separately and then synthesized to provide a broader view of the problem. As stated in Ponce and Pagán Maldonado 2015 [39], “The quantitative approach measures the objective aspects of the problem and the qualitative phase enters the subjective aspects of the problem or the experiences of the participants.” A consistency analysis was also completed between the results of the qualitative and quantitative data collected for three quantitative items to establish validity.
Since the survey did not directly assess student ability, only their perceptions of their ability, we cannot make strong claims about changes in their actual abilities or skills, only their confidence. Post-intervention surveys also included direct questions about their experience and the quality of the assignment itself. Student responses were not anonymous, but identifying information was only used to match pre-survey and post-survey responses and categorize students to their instructor. Additionally, instructors participated in post-lesson interviews to gather qualitative information about the efficacy of the interventions and avenues for future improvements. All study protocols and assessments described within were approved by both UCLA and University of Virginia’s Institutional Review Boards (UCLA IRB#22-001473 and UVA IRB#5514). Full text of the assessment questions is available in Appendix B.
We measured the internal consistency of our assessment using Cronbach’s alpha, which tells us the degree to which the questions in each of the six conceptual categories are closely related. This statistic is commonly used in physics education research [40, 41, 42]. An acceptable value for Cronbach’s alpha is , with a value considered excellent [43]. We calculated Cronbach’s alpha using the same dataset from 52 students as used in the remainder of the analysis, described further in the following section. Table 1 gives the Cronbach’s alpha values for each conceptual category, as calculated for the pre-survey and post-survey responses separately. As the Cronbach’s alpha values all fell within acceptable limits, we felt comfortable proceeding with this model of the conceptual categories for our analysis.
| Category | Question numbers in category | Pre-Survey () | Post-Survey () |
| Jargon | 1-5 | 0.776 | 0.743 |
| Main Takeaways | 6-11 | 0.915 | 0.878 |
| Conceptual | 12-18 | 0.925 | 0.885 |
| Communication | 19-24 | 0.937 | 0.928 |
| Ability | 25-29 | 0.888 | 0.857 |
| Belonging | 30-35 | 0.906 | 0.897 |
II.3 Recruitment & Study Population
The data used in this study is a subset of a larger dataset including all four Astrobites-related assignments. Although the whole dataset includes 203 unique students, our sample focuses on the reading assignment only and contains 52 students from seven courses for astronomy-related majors taught by six instructors across six institutions from Fall 2022 to Winter/Spring 2023 who completed both the pre-survey and the post-survey. Student participants were undergraduates with an intention to major in astronomy or a related subject, and were either taking the course featuring the assignment from Astrobites for major or minor credit, or intending to pursue an astronomy-related career. Five of the host institutions (hosting six of the courses) were R1 universities (doctoral-granting institutions with a high level of research activity), while one was a liberal arts college (hosting one of the courses). We also note that our sample size is small in part due to the fact that undergraduate astronomy majors are a small population – 2023 AIP reports list enrolled juniors and seniors in physics, but only enrolled in astronomy [44, 45].
Participating instructors were recruited via a public call for participants, posted online on the Astrobites website and emailed to faculty and institutional email lists. No compensation was offered. Instructors were allowed to choose from any of the available Astrobites assignment templates, although the data analyzed in this work only includes those who chose the reading assignment. All instructors were required to participate in a pre-intervention meeting to ensure knowledge of study protocols (including how to distribute student survey assessments). Also required was a post-intervention meeting to gather information on the instructors’ experiences incorporating the interventions in their classes. Each participating instructor corresponded with the research team to implement the assignment in their classroom; research team members only advised the instructor and answered questions, but did not have any direct contact with the students themselves. Students were notified of the surveys to fill out by their instructor, using the instructor’s preferred course communication platform. Instructors notified the study team of any deviations from the assignment template as provided. Some instructors were allowed to repeat an assignment multiple times and relevant data was collected, although the analysis of dosing effects in those data [46] from exposure to multiple lessons is not included in this work and we do not include students’ repeated exposures in this dataset.
III Quantitative Analysis & Results
The data set for quantitative analysis consists of 52 complete pre-survey and post-survey responses to the 35 Likert-scale questions listed in Appendix B. We summarize the change in the responses to each question with the 95% confidence interval (‘CI95%’) for the mean of the pre-survey/post-survey differences in student responses. Using the statistics package Pingouin [47], we perform a paired t-test for each question by pairing pre-survey and post-survey responses for individual students. For each question, the p-value was determined using the canonical significance threshold of 0.05 [48].
The power is the probability of correctly rejecting a false null hypothesis ‘H0’. In this study, higher power (values closer to 1) indicates a lower risk of incorrectly finding no change between the pre-lesson and post-lesson surveys. The Bayes factor in favor of the alternative hypothesis (‘BF10’) (e.g., [49]) is computed and expresses the ratio of the likelihood of the data under the alternative hypothesis, ‘H1’: that student perceptions changed, to the null hypothesis, ‘H0’: that student perceptions did not change. A BF10 value of 1 demonstrates equal evidence for the null and alternative hypotheses. Values above 1 up to 3 constitute anecdotal evidence, from 3 to 10 show moderate evidence, from 10 to 100 show strong to very strong evidence, and values over 100 are considered extreme, as defined in the Lee and Wagenmakers [50] scale. BF10 values less than 1 represent (varying) levels of evidence for the null hypothesis, with smaller values representing stronger evidence for H0 – that is, evidence that there was no effect.
We also analyzed changes within the six designed categories; the same statistical analyses performed with responses to individual questions were applied to each of the conceptual categories. “Reversed” Likert scale questions were adjusted222For quantitative questions 2 and 5, low Likert scores (disagreement) represent greater confidence. This is the opposite of the majority of the questions, for which higher Likert scores indicate greater confidence. by reversing the Likert scale for these responses prior to performing any group calculations. Summaries from the quantitative analysis are shown in Table 2 and Figure 1 for the conceptual categories.
| Group | p-value | CI95% | BF10 | BF10 | power |
| interpretation | |||||
| Jargon | [0.10 0.39] | 15.1 | strong | 0.70203 | |
| Main Takeaways | [0.25 0.51] | extreme | 0.99923 | ||
| Conceptual | [0.05 0.29] | 4.9 | moderate | 0.69154 | |
| Communication | 0.035 | [0.01 0.24] | 0.6 | anecdotal for H0 | 0.36436 |
| Ability | [0.06 0.36] | 2.5 | anecdotal | 0.67631 | |
| Belonging | 0.075 | [0.01 0.25] | 0.3 | anecdotal for H0 | 0.29360 |
Almost all questions showed positive mean of differences, indicating an increase in Likert score, corresponding to an average increase in perceived abilities between the pre-survey and post-survey. The exceptions are questions 14 and 17 (“I am able to relate different astronomy concepts to each other.” and “I am able to draw analogies between concepts discussed in astronomy literature and concepts prevalent in everyday life.”), which showed a mean difference of 0, and questions 26 and 34 (“I am capable of succeeding in my astronomy courses.” and “I have a strong sense of belonging to the community of astronomers.”), which each showed an insignificant () decrease (0.01 points each) in Likert score.